OpenStudy (anonymous):

a car accelerates from rest for a distance \[\alpha s \] , after which it decelerates for a distance \[\beta s\] and comes to rest in total time t..find the maximum velocity and initial acceleration?

OpenStudy (anonymous):

v max =2s \[(\alpha + \beta)/t\] a1=\[2s(\alpha+\beta)^{2}/\alpha t ^{2}\]

OpenStudy (anonymous):

this is the answer

OpenStudy (anonymous):

apply s=ut+1/2a*t^2 for the first interval we have t=t1 & u=0 & s=alpha S hence alpha s=0+1/2*a1*t1^2 also for this interval v=u+a t1 or v=0+a*t1 for the next interval we have beta s=at1*(t-t1)+1/2*a2*(t-t1)^2 & 0=at1+a2*(t-t1) these are the set of equations to solve? can u solve it? i am solving right now..

OpenStudy (anonymous):

i will try to solve

OpenStudy (anonymous):

did nt understand this beta s=at1*(t-t1)+1/2*a2*(t-t1)^2

OpenStudy (anonymous):

yeah i reached the first reult..

OpenStudy (anonymous):

ok plzzz go on

OpenStudy (anonymous):

 from this graph we have area of 1st trinagle \[\alpha s=1/2*t1*V_{\max}\] area of 2nd triangle \[\beta s=1/2*V_{\max}*(t-t1)\] we solve this two equation from ist we have Vmax=2*alpha s/t1 from 2nd we have Vmax=2*beta s/(t-t1) we equate the above two equations to find the value of t1 as t1=alpha t/(alpha + beta) put this in equation for Vmax in 1st equation we get Vmax as desired.